OpenStudy (anonymous):

An equasion of the line tangent to the graph of f(x)=x(1-2x)^3 at the point is? A) y=-7x+6 B) y=-6x+5 C) y=-2x+1 D) y=2x+3 E) y=7x-8

OpenStudy (raden):

what point is it ?

OpenStudy (anonymous):

oops sorry. At the point (1,-1)

OpenStudy (raden):

first, determine the slope (m) by take the first derivative of f(x)

OpenStudy (raden):

use the product rule, and also the chain rule

OpenStudy (anonymous):

(1−2x)^3−6x(1−2x)^2 ?

OpenStudy (raden):

yup, good job now to get m, put x=1 to f ' m = (1−2(1))^3−6(1)(1−2(1))^2 = .... ?

OpenStudy (anonymous):

so the slope is -7. A?

OpenStudy (raden):

i got like that too :) now, use to form tangen line use the formula a straight line equation : y - y1 = m(x - x1) here, given (x1,y1) = (1,-1) and m we got = -7

OpenStudy (anonymous):

y+1=-7(x+1) y=-7x+6 OK thank you!

OpenStudy (raden):

yes, that's right you're welcome